# Generative Active Adaptation for Drifting and Imbalanced Network Intrusion Detection

### v1

@@ 2.4. Motivation
We derive the motivation for NetGuard by evaluating how existing active learning approaches perform on CIC-IDS 2017/2018. Fig. 1 illustrates the latent space distribution of benign and attack classes across the two datasets. In Fig. 1(a), the left graph shows the class distribution in CIC-IDS 2017, while the right graph represents the same classes in CIC-IDS 2018, focusing on common attacks. As evident, while the MLP decision boundaries trained on 2017 CIC data (left graph) remains unchanged (right graph), there is a significant class distribution between the two datasets, confirming the presence of concept drift. The figure demonstrates that the same attack classes occupy different regions in the feature space across datasets, indicating distributional changes between two domains. Similarly, Fig. 1(b) shows DDoS HTTP latent space distribution in CIC-IDS 2017 (left graph), while new DDoS variants emerge in CIC-IDS 2018 (right right). Table 2 presents the sample strength of different attack classes in CIC-IDS 2017 and 2018, highlighting severe class imbalance. The Earth Mover’s Distance (EMD) or Wassertein-1 distance column measures the distance between two probability distributions between the same classes in the two datasets, with higher EMD indicating significant drift. When we apply different AL sampling approaches on CIC-IDS 2018, Uncertainty-based AL (Bilgic et al., 2010) fails to sample from minority classes like Bot, FTP-BruteForce, and Web Attack. Similarly, Uncertainty + Diversity (CLUE) (Prabhu et al., 2021) improves sample selection slightly but still fails to prioritize rare classes. On the other hand, our NetGuard’s density-aware sample selection strategy selects more samples from classes with high EMD (such as Infilteration, DDoS, DoS Slowloris), ensuring that samples representative of evolving distribution shifts are included in retraining. However, AL alone does not resolve class imbalance, as underrepresented attacks remain scarce even with optimized sampling. To address this, we integrate AL with state-of-the-art deep generative models, enabling data augmentation to improve detection of rare attacks.
Class CIC 2017 Samples Strength CIC 2018 Samples Strength Norm. EMD Uncertainty Diversity (CoreSet) Uncertainty + Diversity (CLUE) NetGuard AL (Ours) Benign 37937 23494 1.0000 63 366 159 249 Bot 1956 9682 0.3691 0 24 36 19 DDoS 11555 11018 0.5170 44 57 66 129 DoS GoldenEye 9078 6850 0.1216 578 53 100 30 DoS Hulk 12131 10112 0.1066 5 14 102 26 DoS Slowhttptest 5499 9038 0.2159 0 21 65 53 DoS slowloris 5796 1813 0.2478 37 40 29 61 FTP-BruteForce 7935 9329 0.0616 0 19 50 15 Infiltration 36 9140 0.6161 21 165 61 153 SSH-Bruteforce 5897 9302 0.2062 15 36 131 51 Web Attack 2180 151 0.1073 36 4 0 26
Table 2. Class distributions for CIC datasets, normalized EMD for same classes, no. of samples selected per sampling approach
@@ 3.1. Design Overview
Figure 2 illustrates the end-to-end workflow of NetGuard ,which consists of two key modules: density-aware active learning and data augmentation for minority attack classes. The process begins with training a deep learning-based classifier on the original labeled training dataset ($X_{L}$). Once trained, the classifier is tested on a new, unlabeled testing dataset ($X_{UL}$), where it predicts labels for the unseen data. If the performance metrics degrade, the active learning module is triggered to enhance model performance. In the density-aware active learning stage, a Gaussian Mixture Model (GMM) is fitted to both $X_{L}$ and $X_{UL}$. Using the learned distributional information and an informativeness metric, the most representative and informed samples are selected within a given budget. These selected samples are then queried by a human annotator or an oracle for labeling. The annotator continuously expands the original training dataset by adding the newly labeled samples, which helps improve the model’s generalization capability. Additionally, this step addresses the issue of dataset diversity, ensuring that low-distribution (minority) attack classes are well represented. To further enhance model robustness, minority attack classes with a distribution of less than five percent are fed into the data augmentation module. This module generates synthetic samples to strengthen the presence of these underrepresented attack classes. The newly augmented samples are then incorporated into the training corpus, ensuring a more balanced dataset. By integrating active learning with data augmentation, NetGuard effectively addresses concept drift by selecting the most informative samples for labeling while mitigating imbalanced dataset issues. This iterative learning framework improves the model’s ability to detect evolving and unknown attacks in network intrusion detection systems.
Figure 2. Continuous learning framework of NetGuard
@@ 3.1.1. Problem Formulation
Therefore to solve this problem and given the distributional discrepancy between $X_{L}$ and $X_{UL}$, we devise a new informativeness metric $I_{Score}$ within the active learning framework by drawing inspiration from the well-known unsupervised clustering model, Gaussian Mixture Model (GMM), a powerful way to probabilistically represent latent distribution with its parameter set formalized as $\psi\sim(\pi,\mu,\sigma^{2})$, where $\psi$ encapsulates the mixture weights ($\pi$) component means ($\mu$), and covariances ($\sigma^{2}$) of the GMM. These parameters are learned using the Expectation-Maximization (EM) algorithm. Using the learned distributional information allows us to estimate the log-likelihood ($p(z)$) of unlabeled samples $x_{UL}$ with respect to both the labeled dataset $X_{L}$ and the unlabeled dataset $X_{UL}$ which is then embedded into the selection criterion to compute $I_{Score}$. Samples with high $I_{Score}$ values are prioritized within the constrained labeling budget $B$ for retraining the model.
@@ 3.1.2. Workflow
In each iteration, the classifier selects the most informative samples from unlabeled data $X_{UL}$ based on a predefined informativeness criterion, queries an oracle for labels, and expands $X_{L}$. This process continues until the labeling budget $B$ is exhausted or a stopping condition is met.
  1. (1)
    Initial Model Training 1 In AL, a classifier $f$ is trained in iterative rounds $t=0,1,\dots,T$, starting with a DNN based model initially trained on the labeled dataset $X_{L}$.
  2. (2)
@@
    Fit GMMs to $X_{L}$ and $X_{UL}$ 3 4 To characterize the latent distributions of the labeled and unlabeled datasets, we fit a GMM to the labeled dataset $X_{L}$ to obtain the parameters $\psi_{L}=(\pi_{L},\mu_{L},\sigma_{L}^{2})$. Similarly, another GMM is fitted to the unlabeled dataset $X_{UL}$ to obtain the parameters
    $\psi_{UL}=(\pi_{UL},\mu_{UL},\sigma_{UL}^{2})$.
  4. (4)
@@
    Compute Log-Likelihoods 5 The intuition behind sample selection is that those samples should be chosen which are dissimilar to $X_{L}$ but at the same time they are best representatives of $X_{UL}$. Subseqeuntly, using the learned GMMs, log-likelihood is computed for each unlabeled sample $x_{UL}$. First, we compute its log-likelihood under the GMM of the unlabeled dataset, resulting in:
    $p(z_{UL};\psi_{UL})=\sum_{j=1}^{K}\pi_{UL,j}\mathcal{N}(z_{UL}|\mu_{UL,j},\sigma_{UL,j}^{2})$
    Then, the log-likelihood of the same sample under the GMM of the labeled dataset is computed as:
    $p(z_{UL};\psi_{L})=\sum_{j=1}^{K}\pi_{L,j}\mathcal{N}(z_{UL}|\mu_{L,j},\sigma_{L,j}^{2})$
    Note that the log-likelihood $p(z_{UL};\psi_{UL})$ is calculated as it represents the probability that a given data of interest is generated by the GMM characterized by $\psi$.
  5. (5)
    Calculate Informativeness Score 5 The informativeness score for a data point $x_{UL}$ is defined as the difference between its log-likelihood under the unlabeled dataset’s GMMand its log-likelihood under the labeled dataset’s GMM. Mathematically, this is expressed as:
    (1) $I_{\text{Score}}(x_{UL})=p(x_{UL};\psi_{UL})-p(x_{UL};\psi_{L})$
    This score quantifies the most valuable data points $x_{UL}$ as they help the model reduce distributional gaps between the training and testing datasets and improve generalization.
  6. (6)
    Sample Selection Strategy 5 The acquisition function $A(x)$ is defined as the set of all informativeness scores $I_{\text{score}}$ values for $x\in X_{UL}$:
    (2) $A(x)=\{I_{\text{score}}(x)\mid x\in X_{UL}\}$
@@
    To select the most informative samples, the scores are sorted in descending order, and the top $B$ samples are chosen based on the available labeling budget $B$. This ensures that the batch $X_{B}\subseteq X_{UL}$ consists of the most informative samples for labeling. It is worth mentioning that by sorting samples in descending order of their informativeness score, we prioritize those with high $I_{\text{Score}}$ thereby selecting samples that both improve coverage and maximize informativeness. Samples that improve coverage are those with high $p(x_{UL};\psi_{UL})$ that are representative of the unlabeled dataset and help the model learn from regions of the feature space that are underrepresented in $X_{L}$. Meanwhile, samples that maximize informativeness are those with low $p(x_{UL};\psi_{L})$ which are essentially dissimilar to the labeled dataset $X_{L}$ and likely come from regions where the model’s predictions are uncertain.
  7. (7)
@@
    Obtain Ground Truth Labels 6 Send the selected subset $X_{B}$ to the human annotator (or an oracle) to obtain ground-truth labels.
  8. (8)
@@
    Update Labeled/Unlabeled Sets 7 Add these newly labeled samples $(X_{B},Y_{B})$ to the labeled training set $X_{L}$. Remove $X_{B}$ from unlabeled dataset $X_{UL}$.
  9. (9)
    Train/Re-Train Model 8 Incorporate the updated labeled set $X_{L}$ into the training process to improve the model’s performance.
3.2. Diversity-driven Data Augmentation
@@
After the active learning step, the sampled points are included in the original labeled dataset ($X_{L}$), which still remains imbalanced with under-represented classes, referred to as the minority classes. A threshold of 5% is used to identify minority classes. The goal is to synthesize new samples using deep generative techniques 8 such that (a) the sample count of each minority class matches that of the majority class, ensuring class balance, (b) the generated samples maintain representativeness within a single class by capturing the underlying distribution of each minority class, but they should be different enough to prevent redundancy (near duplicate samples) and subsequent model overfitting and (c) the synthetic samples generated for different classes should remain distinct from each other with clear separation between classes in feature space.
3.2.1. Problem Formulation.
@@
Let $X_{B}$ be the set of samples obtained from the density-aware active learning module using a given labeling budget $B$. After being labeled by the Oracle, each sample in $X_{B}$ corresponds to a row in the tabular dataset, where each row is a $d$-dimensional feature vector. The corresponding class labels are denoted as $Y_{B}$. Assume this batch contains $C_{B}$ different classes. These actively learned labeled samples $(X_{B},Y_{B})$ are added to the existing labeled training dataset $X_{L}$, forming an updated dataset $X^{\prime}_{L}$ with labels $Y^{\prime}_{L}$. The expanded dataset spans a set of classes $C^{\prime}$, which includes previously seen classes from $X_{L}$ and any newly introduced classes from $X_{B}$. Despite the expansion, the dataset may still be imbalanced, with some classes having fewer than 5% of the total samples. Let $C^{\prime}_{\min}$ represent the set of such minority classes, and let $X^{\prime}_{\text{min}}$ be the subset of training samples that belong to these minority classes, $X^{\prime}_{\text{min}}=\{(x^{\prime}_{i},y^{\prime}_{i})\mid y^{\prime}_{i}\in C^{\prime}_{\min}\}$. Let $G$ be a generative model trained on the minority classes $X^{\prime}_{\text{min}}$.
3.2.2. Deep Generative Training.
@@
The goal of $G$ is to learn the true joint distribution on the different metadata and measurement features of minority classes, represented as $P_{\text{min}}(x^{\prime})$. The model $G$, parameterized by $\theta$, learns an approximation $P_{Gen}(x^{\prime};\theta)$ on the true distribution. $G$ is trained by minimizing the loss function $L_{G}$ that measures the divergence between the actual minority class distribution $P_{\text{min}}(x^{\prime})$ and the generated distribution $P_{Gen}(x^{\prime};\theta)$:
$\theta^{*}=\arg\min_{\theta}L_{G}(P_{\text{min}}(x^{\prime}),P_{Gen}(x^{\prime};\theta))$
where $L_{G}$ is the divergence measure such as Kullback-Leibler (KL) divergence. Once trained, $G$ generates synthetic samples by sampling from different regions of its learned distribution.
@@
For example, consider a minority class where the metadata feature $m_{1}$ (protocol type) has states such as $s_{1}$ (HTTP), $s_{2}$ (FTP), $s_{3}$ (SSH) and the metadata feature $m_{2}$ (source IP range) has states $s_{4}$ (range A), $s_{5}$ (range B). The generative model can interpolate between these states, such as generating samples with combinations like $(s_{1},s_{5}),(s_{2},s_{7})$ or $(s_{3},s_{9})$ while conditioning on measurement features like $z_{1}$ (packet size) and $z_{2}$ (flow duration). Formally, we can represent it as $x^{\prime}_{\text{Gen}}=G\bigl{(}s_{i},s_{j}\;\big{|}\;z_{1},z_{2}\bigr{)}$, where $x^{\prime}_{\text{Gen}}$ is a synthetic sample generated from the minority class by combining metadata states $s_{i},s_{j}$ with measurement features $z_{1},z_{2}$. This interpolation creates variations in the synthetic samples, ensuring that the generated data captures a wide range of scenarios within the minority class. In practice, the choice of generative models implicitly address diversity in synthetic generation by their distinct mechanisms for learning and sampling from the joint distribution of features. The augmented dataset is then formed by combining these synthetic samples with the original training data, improving class balance while preserving diversity.
3.2.3. Post-Processing.
After generating synthetic samples, a post-processing step is applied to remove low-quality samples that may resemble benign data. This is done using a binary logistic regression classifier trained to distinguish between benign and non-benign samples. The classifier is trained on the original labeled dataset $X_{L}$ and then used to evaluate the generated synthetic samples. Since all synthetic samples belong to the non-benign class, only those with a low probability of being benign are retained. This filtering step ensures that synthetic samples do not closely resemble majority-class benign samples, improving the overall quality of the augmented dataset. By applying this post-processing step, we effectively clean the synthetic data and remove samples that may not contribute to improving minority class representation.
3.2.4. Validation and Classification.
We evaluate the fidelity of synthetic samples using two approaches: (i) by computing the Wasserstein-2 (W2) distance to ensure the overall structure and spacing of synthetic data points are preserved, and (ii) by assessing the performance of downstream classification tasks.
To achieve this, after the augmentation step, the final training set is constructed by combining all minority samples (both synthetic and original) with the majority class samples from the original training dataset $X_{L}$ as well as the actively learned samples $X_{B}$. We train both the classifiers on the expanded dataset. For evaluation, we use the future testing dataset $X_{UL}$ such that the samples used during the active learning phase i.e. $X_{B}$ are excluded from $X_{UL}$.

### v2

@@ 2.2. Motivation
We derive the motivation for NetGuard by evaluating how existing active adaptation approaches perform on CIC-IDS 2017/2018. Fig. 1 illustrates the latent space distribution of benign and attack classes across the two datasets. In Fig. 1(a), the left graph shows the class distribution in CIC-IDS 2017, while the right graph represents the same classes in CIC-IDS 2018, focusing on common attacks. As evident, while the MLP decision boundaries trained on 2017 CIC data (left graph) remain unchanged (right graph), there is a significant class distribution between the two datasets, confirming the presence of concept drift. The figure demonstrates that the same attack classes occupy different regions in the feature space across datasets, indicating distributional changes between two domains. Similarly, Fig. 1(b) shows DDoS HTTP latent space distribution in CIC-IDS 2017 (left graph), while new DDoS variants emerge in CIC-IDS 2018 (right right). Table 2 presents the sample strength of different attack classes in CIC-IDS 2017 and 2018, highlighting severe class imbalance. The Earth Mover’s Distance (EMD) or Wassertein-1 distance column measures the distance between two probability distributions between the same classes in the two datasets, with higher EMD indicating significant drift. When we apply different AL sampling approaches on CIC-IDS 2018, Uncertainty-based AL (Bilgic et al., 2010) fails to sample from minority classes like Bot, FTP-BruteForce, and Web Attack. Similarly, Uncertainty + Diversity (CLUE) (Prabhu et al., 2021) improves sample selection slightly but still fails to prioritize rare classes. On the other hand, our NetGuard’s density-aware sample selection strategy selects more samples from classes with high EMD (such as Infiltration, DDoS, DoS Slowloris), ensuring that samples representative of evolving distribution shifts are included in retraining. However, it alone does not resolve class imbalance, as underrepresented attacks remain scarce even with optimized sampling. To address this, we adapt it as an approach to select priors for state-of-the-art deep generative models, enabling data augmentation to improve the detection of rare attacks.
Class CIC 2017 #Samples CIC 2018 #Samples Norm. EMD Uncertainty Diversity (CoreSet) Uncert. + Diversity (CLUE) NetGuard PS Benign 37937 23494 1.0000 63 366 159 249 Bot 1956 9682 0.3691 0 24 36 19 DDoS 11555 11018 0.5170 44 57 66 129 DoS GoldenEye 9078 6850 0.1216 578 53 100 30 DoS Hulk 12131 10112 0.1066 5 14 102 26 Slowhttptest 5499 9038 0.2159 0 21 65 53 Slowloris 5796 1813 0.2478 37 40 29 61 FTP-BruteForce 7935 9329 0.0616 0 19 50 15 Infiltration 36 9140 0.6161 21 165 61 153 SSH-BruteForce 5897 9302 0.2062 15 36 131 51 Web Attack 2180 151 0.1073 36 4 0 26
Table 2. Class distributions for CIC datasets, normalized EMD, and samples selected per strategy. PS = Prior Selection.
@@ 3.1. Design Overview
Figure 2 illustrates the end-to-end workflow of NetGuard, which consists of two key modules: density-aware dataset prior selection and conditional data augmentation for minority attack classes. The process begins with training a deep learning-based classifier on the original labeled training dataset ($X_{L}$). Once trained, the classifier is tested on a new, unlabeled testing dataset ($X_{UL}$), where it predicts labels for the unseen data. If the performance metrics degrade, the prior selection module is triggered to enhance model performance. In the density-aware dataset prior selection stage, a Gaussian Mixture Model (GMM) is fitted to both $X_{L}$ and $X_{UL}$. Using the learned distributional information and an informativeness metric, the most representative and informed samples $X_{P}$ are selected within a given budget. These selected samples are then queried by a human annotator or an oracle for labeling. The annotator continuously expands the original training dataset by adding the newly labeled samples, which helps improve the model’s generalization capability. Additionally, this step addresses the issue of dataset diversity, ensuring that low-distribution (minority) attack classes are well represented. To further enhance model robustness, minority attack classes with less than 5% distribution are fed into the data augmentation module. Unlike generic augmentation, it performs conditional generation synthesizing samples conditioned on the actively selected, high-informative priors from $X_{UL}$. This ensures that synthetic samples reflect the current test-time distribution and preserve attack semantics. The generated samples are integrated into the training corpus, producing a more balanced dataset and strengthening the model’s robustness to detect underrepresented and evolving attacks. Thus, NetGuard introduces a novel conditional data augmentation strategy, anchored on actively selected samples, to improve detection under concept drift and class imbalance.
Figure 2. Closed-loop learning framework of NetGuard
3.2. Density-aware Dataset Prior Selection
@@ 3.2.1. Problem Formulation
Therefore, to solve this problem and given the distributional discrepancy between $X_{L}$ and $X_{UL}$, we devise a new informativeness metric $I_{Score}$ within the prior selection framework by drawing inspiration from the well-known unsupervised clustering model, Gaussian Mixture Model (GMM), a powerful way to probabilistically represent latent distribution with its parameter set formalized as $\psi\sim(\pi,\mu,\sigma^{2})$, where $\psi$ encapsulates the mixture weights ($\pi$) component means ($\mu$), and covariances ($\sigma^{2}$) of the GMM. These parameters are learned using the Expectation-Maximization (EM) algorithm. Using the learned distributional information allows us to estimate the log-likelihood ($p(z)$) of unlabeled samples $x_{UL}$ with respect to both the labeled dataset $X_{L}$ and the unlabeled dataset $X_{UL}$ which is then embedded into the selection criterion to compute $I_{Score}$. Samples with high $I_{Score}$ values are prioritized within the Prior labeling budget $P$ for retraining the model.
@@ 3.2.2. Workflow
In each iteration, the classifier selects the most informative samples from unlabeled data $X_{UL}$ based on a predefined informativeness criterion, queries an oracle for labels, and expands $X_{L}$. This process continues until the labeling budget $P$ is exhausted or a stopping condition is met.
Initial Model Training 1 In AL, a classifier $f$ is trained in iterative rounds $t=0,1,\dots,T$, starting with a DNN-based model initially trained on the labeled dataset $X_{L}$.
@@
Fit GMMs to $X_{L}$ and $X_{UL}$ 3 4 To characterize the latent distributions of the labeled and unlabeled datasets, we fit a GMM to the labeled dataset $X_{L}$ to obtain the parameters $\psi_{L}=(\pi_{L},\mu_{L},\sigma_{L}^{2})$. Similarly, another GMM is fitted to the unlabeled dataset $X_{UL}$ to obtain the parameters $\psi_{UL}=(\pi_{UL},\mu_{UL},\sigma_{UL}^{2})$.
@@
Compute Log-Likelihoods 5 The intuition behind sample selection is that those samples should be chosen which are dissimilar to $X_{L}$, but at the same time they are the best representatives of $X_{UL}$. Subsequently, using the learned GMMs, the log-likelihood is computed for each unlabeled sample $x_{UL}$. First, we compute its log-likelihood under the GMM of the unlabeled dataset, resulting in: $p(z_{UL};\psi_{UL})=\sum_{j=1}^{K}\pi_{UL,j}\mathcal{N}(z_{UL}|\mu_{UL,j},\sigma_{UL,j}^{2})$. Then, the log-likelihood of the same sample under the GMM of the labeled dataset is computed as: $p(z_{UL};\psi_{L})=\sum_{j=1}^{K}\pi_{L,j}\mathcal{N}(z_{UL}|\mu_{L,j},\sigma_{L,j}^{2})$. Note that the log-likelihood $p(z_{UL};\psi_{UL})$ is calculated as it represents the probability that a given data of interest is generated by the GMM characterized by $\psi$.
@@
Calculate Informativeness Score 5 The informativeness score for a data point $x_{UL}$ is defined as the difference between its log-likelihood under the unlabeled dataset’s GMM and its log-likelihood under the labeled dataset’s GMM. Notationally, this is $I_{\text{Score}}(x_{UL})=p(x_{UL};\psi_{UL})-p(x_{UL};\psi_{L})$. This score quantifies the most valuable data points $x_{UL}$ as they help the model reduce distributional gaps between the training and testing datasets. Prior Selection Strategy 5 The acquisition function $A(x)$ is defined as the set of all informativeness scores $I_{\text{score}}$ values for $x\in X_{UL}$: $A(x)=\{I_{\text{score}}(x)\mid x\in X_{UL}\}$. To select informative samples, the scores are sorted in descending order, and the top $P$ samples are chosen based on the available labeling budget $P$. This ensures that the batch $X_{P}\subseteq X_{UL}$ consists of the most informative samples for labeling. It is worth mentioning that by sorting samples in descending order of their informativeness score, we prioritize those with high $I_{\text{Score}}$, thereby selecting samples that both improve coverage and maximize informativeness. Samples that improve coverage are those with high $p(x_{UL};\psi_{UL})$ that are representative of the unlabeled dataset and help the model learn from regions of the feature space that are underrepresented in $X_{L}$. Meanwhile, samples that maximize informativeness are those with low $p(x_{UL};\psi_{L}),$ which are essentially dissimilar to the labeled dataset $X_{L}$ and likely come from regions where the model’s predictions are uncertain.
@@
Obtain Ground Truth Labels 6 Send the selected subset $X_{P}$ to the human annotator (or an oracle) to obtain ground-truth labels.
@@
Update Labeled/Unlabeled Sets 7 Add these newly labeled samples $(X_{P},Y_{P})$ to the labeled training set $X_{L}$. Remove $X_{P}$ from unlabeled dataset $X_{UL}$.
3.3. Diversity-aware Conditional Data Augmentation
@@
After the prior selection step, the sampled points are included in the original labeled dataset ($X_{L}$), which remains imbalanced with under-represented classes, referred to as the minority classes. A threshold of 5% is used to identify minority classes. The goal is to synthesize new samples using deep generative techniques 8 such that (a) the sample count of each minority class matches that of the majority class, ensuring class balance, (b) the generated samples maintain representativeness within a single class by capturing the underlying distribution of each minority class, but they should be different enough to prevent redundancy (near duplicate samples) and subsequent model overfitting and (c) the synthetic samples generated for different classes should remain distinct from each other with clear separation between classes in feature space.
@@
Problem Formulation. Let $X_{P}$ be the set of samples obtained from the density-aware prior selection module using a given labeling budget $P$. After being labeled by the Oracle, each sample in $X_{P}$ corresponds to a row in the tabular dataset, where each row is a $d$-dimensional feature vector. The corresponding class labels are denoted as $Y_{P}$. Assume this batch contains $C_{P}$ different classes. These actively learned labeled samples $(X_{P},Y_{P})$ are added to the existing labeled training dataset $X_{L}$, forming an updated dataset $X^{\prime}_{L}$ with labels $Y^{\prime}_{L}$. The expanded dataset spans a set of classes $C^{\prime}$, which includes previously seen classes from $X_{L}$ and any newly introduced classes from $X_{P}$. Despite the expansion, the dataset may still be imbalanced, with some classes having fewer than 5% of the total samples. Let $C^{\prime}_{\min}$ represent the set of such minority classes, and let $X^{\prime}_{\text{min}}$ be the subset of training samples that belong to these minority classes, $X^{\prime}_{\text{min}}=\{(x^{\prime}_{i},y^{\prime}_{i})\mid y^{\prime}_{i}\in C^{\prime}_{\min}\}$. Let $G$ be a generative model trained on the minority classes $X^{\prime}_{\text{min}}$.
@@
Deep Generative Training. The goal of $G$ is to learn the true joint distribution on the different metadata and measurement features of minority classes, represented as $P_{\text{min}}(x^{\prime})$. The model $G$, parameterized by $\theta$, learns an approximation $P_{Gen}(x^{\prime};\theta)$ on the true distribution. $G$ is trained by minimizing the loss function $L_{G}$ that measures the divergence between the actual minority class distribution $P_{\text{min}}(x^{\prime})$ and the generated distribution $P_{Gen}(x^{\prime};\theta)$: $\theta^{*}=\arg\min_{\theta}L_{G}(P_{\text{min}}(x^{\prime}),P_{Gen}(x^{\prime};\theta))$, where $L_{G}$ is the divergence measure, such as Kullback-Leibler (KL) divergence. Once trained, $G$ generates synthetic samples by sampling from different regions of its learned distribution.
@@
For example, consider a minority class where the metadata feature $m_{1}$ (protocol type) has states such as $s_{1}$ (HTTP), $s_{2}$ (FTP), $s_{3}$ (SSH) and the metadata feature $m_{2}$ (source IP range) has states $s_{4}$ (range A), $s_{5}$ (range B). The generative model can interpolate between these states, such as generating samples with combinations like $(s_{1},s_{5}),(s_{2},s_{7})$ or $(s_{3},s_{9})$ while conditioning on measurement features like $z_{1}$ (packet size) and $z_{2}$ (flow duration). Formally, we can represent it as $x^{\prime}_{\text{Gen}}=G\bigl{(}s_{i},s_{j}\;\big{|}\;z_{1},z_{2}\bigr{)}$, where $x^{\prime}_{\text{Gen}}$ is a synthetic sample generated from the minority class by combining metadata states $s_{i},s_{j}$ with measurement features $z_{1},z_{2}$. This interpolation creates variations in the synthetic samples, ensuring that the generated data captures a wide range of scenarios within the minority class. In practice, the choice of generative models implicitly addresses diversity in synthetic generation by their distinct mechanisms for learning and sampling from the joint distribution of features. The augmented dataset is then formed by combining these synthetic samples with the original training data, improving class balance while preserving diversity.
Post-Processing. After generating synthetic samples, a post-processing step is applied to remove low-quality samples that may resemble benign data. This is done using a binary logistic regression classifier trained to distinguish between benign and non-benign samples. The classifier is trained on the original labeled dataset $X_{L}$ and then used to evaluate the generated synthetic samples. Since all synthetic samples belong to the non-benign class, only those with a low probability of being benign are retained. This filtering step ensures that synthetic samples do not closely resemble majority-class benign samples, improving the overall quality of the augmented dataset. By applying this post-processing step, we effectively clean the synthetic data and remove samples that may not contribute to improving minority class representation.
Validation and Classification. We evaluate the fidelity of synthetic samples using two approaches: (i) by computing the Wasserstein-2 (W2) distance to ensure the overall structure and spacing of synthetic data points are preserved, and (ii) by assessing the performance of downstream classification tasks.
To achieve this, after the augmentation step, the final training set is constructed by combining all minority samples (both synthetic and original) with the majority class samples from the original training dataset $X_{L}$ as well as the actively selected samples $X_{P}$. We train both the classifiers on the expanded dataset. For evaluation, we use the future testing dataset $X_{UL}$ such that the samples used during the prior selection phase i.e., $X_{P}$ are excluded from $X_{UL}$.